OpenStudy (anonymous):

Please Help & Explain 3/x= (3?)

OpenStudy (anonymous):

is this a fraction or division

OpenStudy (anonymous):

Division

OpenStudy (anonymous):

then x could = either 1 or 9

OpenStudy (anonymous):

why would it have to be one of those?

OpenStudy (anonymous):

well if its division 3 divided by 9 can give you 3 or 3 divided by 1 can also give you

OpenStudy (anonymous):

Full problem is: 4x^2 + 5x + 3 ------------- I know it's 4x + 5 + (?) x

OpenStudy (raden_zaikaria):

3/x

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

would it be 3?

OpenStudy (raden_zaikaria):

cannot....

OpenStudy (anonymous):

@frankcookie96 your original idea was correct: \[\frac{x^2+5x+3}{x}=\frac{x^2}{x}+\frac{5x}{x}+\frac{3}{x}\]

OpenStudy (anonymous):

so as the result it would be?

OpenStudy (raden_zaikaria):

because 3 cannot divide by x or any unknown....

OpenStudy (anonymous):

Well you simplify your first two terms and get what?

OpenStudy (anonymous):

the 1st one is \[4x ^{2}\] over x which is 4x

OpenStudy (anonymous):

@raden_zaikaria the problem is simplifying an expression without knowing x. You can in fact, divide by x.

OpenStudy (anonymous):

For the 2nd 5x/x I get 5

OpenStudy (anonymous):

So it's 4x + 5 + ?

OpenStudy (anonymous):

I copied the equation wrong, apologies. Yes your first term becomes 4x and your second term is 5. Put all three terms together and you get: \[4x+5+\frac{3}{x}\]

OpenStudy (anonymous):

3/x does not simplify any further.

OpenStudy (anonymous):

Ok thanks! I understand. =)

OpenStudy (anonymous):

Perfect! Thank you for the medal. :)

OpenStudy (anonymous):

Glad to.